OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 2) = 4x. negative 2 over 5 negative 5 over 2 2 over 5 5

OpenStudy (anonymous):

well solve for x to get rid of the fraction first multiply through by 3

OpenStudy (anonymous):

That gives you 2(x-2) = 12x Next distribute 2x - 4 = 12x can you solve for x from this point

OpenStudy (anonymous):

Wouldnt you add 2x and 12x

OpenStudy (anonymous):

not quite - they are on different sides of the equation so you would need to subtract 2x from both sides ok

OpenStudy (anonymous):

So -4=10x

OpenStudy (anonymous):

yes thats it :) Now what would you do to solve for x

OpenStudy (anonymous):

From then im not really sure....sorry i really didnt understand this thats why i posted for help :(

OpenStudy (anonymous):

well we need x by itself. Since we have -4 = 10x We would divide both sides by 10 which give you -4/10 = x and then you reduce your answer which give you x = -2/5 ok:)

OpenStudy (anonymous):

Thanks(: Soo much!

OpenStudy (anonymous):

you are welcome